OpenStudy (anonymous):

Part A: Angelina plans to put a fence along the length AD of her lawn. What is the length of the fence required? Part B: Using complete sentences, explain how you arrived at the answer for Part A.

OpenStudy (nottim):

there a diagram

OpenStudy (nottim):

use pythagorean theorem to figure out BD first.

OpenStudy (zepp):

Hint: Pythagorean Theorem and the 30-60-90 theorem http://puu.sh/qsq6 http://www.themathpage.com/atrig/30-60-90-triangle.htm

OpenStudy (nottim):

aww.. to fast for me zepp

OpenStudy (zepp):

Or, law of sines.

OpenStudy (anonymous):

what?

OpenStudy (nottim):

website seems more complex than necessary for this question tho zepp

OpenStudy (nottim):

u kno what pythagorean theorem is right.

OpenStudy (zepp):

Alright then, forget the 30-60-90 theorem, can you find the measure for BD?

OpenStudy (nottim):

\[c ^{2}=a ^{2}+b ^{?}\]

OpenStudy (nottim):

last ? was a 2

OpenStudy (anonymous):

12 squared + 35 squared= C squared?

OpenStudy (nottim):

yup

OpenStudy (anonymous):

1369 is C?

OpenStudy (nottim):

now sqrt that.

OpenStudy (zepp):

2 squared + 35 squared= C^2 1369 = C^2 You forgot a step ;)

OpenStudy (nottim):

yup

OpenStudy (anonymous):

1,874,161

OpenStudy (zepp):

That's 1369 squared If you want C and you have 1369 = C^2 You have to do \[\sqrt(1369) = \sqrt(C^2)\] \[\sqrt(1369) = C\]

OpenStudy (anonymous):

37

OpenStudy (zepp):

Alright, now you have the length of BD Find the red colored angle I drawn on this http://puu.sh/qsx0

OpenStudy (anonymous):

30

OpenStudy (zepp):

When you have it, use \[sinA/A = sinB/B\]

OpenStudy (anonymous):

...huh?

OpenStudy (zepp):

Do you know the law of sines?

OpenStudy (anonymous):

No I don't think so

OpenStudy (zepp):

Alright, you have to use the 30 - 60 - 90 theorem, look at this http://puu.sh/qsAi This means that the short leg / long leg 's ratio is 1: sqrt(3), (If the short leg's length is 1, long's leg's length would be sqrt(3)) short leg / hypotenuse is 1:2 etc etc. In you picture, BD is the length opposing the 60 degree angle, it's the long leg, and what you have to find is AD, which is the short leg By following the ratio, you have short leg / long leg, ratio is 1: sqrt(3) \[1 : \sqrt(3) = x : 37\] \[1/ \sqrt(3) = x/37\] And solve it, x will be the short leg(AD) 's length, good luck :)

OpenStudy (anonymous):

oh my god I don't get it..

OpenStudy (zepp):

Lemme draw something hold on

OpenStudy (zepp):

http://puu.sh/qsI5 Do you understand this ratio?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I got like 0.15

OpenStudy (zepp):

Alright, what we need in this situation is the length of AD, we already know the length of BD, which is 37ft, we will do this: http://puu.sh/qsK0

OpenStudy (zepp):

Oops, sorry, it's this one: http://puu.sh/qsKI

OpenStudy (anonymous):

okayy

OpenStudy (zepp):

Basically, what this is saying, is that in a rectangle triangle, if the side facing the 60 degree angle is the long leg, if the short leg = 1, the long leg would be = sqrt(3), and based on this, you'll find the solution to your problem

OpenStudy (anonymous):

okay! Thanks!

OpenStudy (zepp):

So you'll get \[\frac{1}{\sqrt{3}} = \frac{x}{37}\] x = 21.36 You're welcome :)